# Counterfactual Tuning for Temporal Sensitivity Enhancement in Large Language Model-based Recommendation

### v1

@@ Abstract.
To address this critical gap, we propose Counterfactual Enhanced Temporal Framework for LLM‑Based Recommendation (CETRec). CETRec is grounded in causal inference principles, which allow it to isolate and measure the specific impact of temporal information on recommendation outcomes. By conceptualizing temporal order as an independent causal factor distinct from item content, we can quantify its unique contribution through counterfactual reasoning—comparing what recommendations would be made with and without temporal information while keeping all other factors constant. This causal framing enables CETRec to design a novel counterfactual tuning objective that directly optimizes the model’s temporal sensitivity, teaching LLMs to recognize both absolute timestamps and relative ordering patterns in user histories. Combined with our counterfactual tuning task derived from causal analysis, CETRec effectively enhances LLMs’ awareness of both absolute order (how recently items were interacted with) and relative order (the sequential relationships between items). Extensive experiments on real-world datasets demonstrate that CETRec consistently outperforms state-of-the-art baselines.
@@ 1. Introduction
In the era of information explosion, recommendation systems have become essential tools that help users navigate vast amounts of content across online platforms (Fan et al., 2023). Within this domain, sequential recommendation has gained increasing attention for its ability to effectively model users’ temporal interaction sequences and capture evolving preferences over time, producing more personalized and accurate recommendations (Wu et al., 2023; Tang and Wang, 2018; Hidasi et al., 2016; Kang and McAuley, 2018). Recently, with the rapid advancement of large language models (LLMs), numerous studies have begun to incorporate these powerful models into sequential recommendation systems, leveraging their rich pre-trained knowledge and reasoning capabilities (Bao et al., 2023a; Zheng et al., 2024; Liao et al., 2024; Geng et al., 2022; Bao et al., 2023b; Wei et al., 2024). Clearly, temporal information—the chronological ordering and timing of user interactions—plays a crucial role in sequential recommendation, as it directly reflects the evolution of user interest and preference dynamics.
@@
Existing LLM-based recommendation approaches can be largely grouped into two main streams: 1) Paradigm adaptation methods that attempt to reformulate recommendation tasks into formats suitable for LLMs’ text understanding and reasoning capabilities (Bao et al., 2023a; Geng et al., 2022). These approaches typically transform recommendation into natural language generation tasks, framing item suggestions as text completion or response generation problems. They leverage prompting techniques and fine-tuning strategies to align LLMs’ pre-trained knowledge with recommendation objectives. 2) Multimodal integration methods treat items as a distinct data modality (Yang et al., 2023; Lin et al., 2023). These approaches construct specialized item embeddings that capture item characteristics, user preferences, and interaction patterns, then integrate these representations with LLMs’ text understanding capabilities through cross-modal attention mechanisms or embedding alignment techniques, leveraging LLMs’ rich pre-trained knowledge to enhance representation of items. Despite these advancements, these approaches primarily focus on modeling the semantic aspects of interacted items or optimizing the overall recommendation task, and do not explicitly capture the temporal structure of user interaction sequences. This temporal oversight leaves a significant gap in their ability to model users’ evolving interests and time-dependent preference shifts. An empirical study (Kim et al., 2025) that perturbs interaction sequences shows that LLM-based recommenders exhibit smaller performance drops under different sequence orders compared to traditional models, further demonstrating their limited ability to capture temporal information.
@@
  • We identify and analyze the insufficient use of temporal information in user interaction histories by existing LLM-based recommenders in the sequential recommendation setting. Through systematic analysis, we reveal how architectural limitations in Transformer models and position embedding schemes designed for natural language processing create fundamental barriers to capturing temporal dynamics in user preference evolution.
@@ 2.1. LLM-based Recommendation
  • Paradigm Adaption. These methods leverage the reasoning capabilities of LLMs by aligning them with the recommendation paradigm—reformulated as natural language generation tasks—through techniques such as fine-tuning (Geng et al., 2022; Bao et al., 2023a, b; Zhang et al., 2024a), prompting (Dai et al., 2023), or in-context learning (ICL) (Hou et al., 2024).
Although a prior study (Kim et al., 2025) has noted that LLMs inadequately exploit the temporal ordering in user interaction sequences—attempting to distill sequence information from conventional models—the majority of existing methods still lack explicit temporal modeling within the LLM itself. To our best knowledge, we are the first to harness LLMs’ native capacity to directly model the temporal order of interaction sequence.
@@ 2.2. Causal Inference for Recommendation
In recommender systems, non‐causal approaches typically model only correlations between variables, which can lead to spurious relationships and limited generalization. In contrast, causal inference aims to reveal the underlying causal mechanisms in the data, enabling more robust and interpretable recommendations (Luo et al., 2023).
Prior work in causal recommendation has primarily focused on three key areas: First, debiasing methods (Zhu et al., 2020; Christakopoulou et al., 2020) employ causal techniques to mitigate various forms of systematic bias in recommendation data. These include popularity bias (Craswell et al., 2008), where frequently interacted items receive disproportionate exposure regardless of actual user preference, and click bias (Wang et al., 2021), where user interaction patterns are influenced by item position rather than genuine interest. By modeling these biases as confounding factors in a causal graph, researchers have developed intervention strategies that produce more balanced and fair recommendations.
Second, causal approaches have proven effective for handling data sparsity and noise (Xiong et al., 2021; Zhang et al., 2021; Liu et al., 2022). By leveraging structural causal models to distinguish genuine user preferences from random interactions or missing data patterns, these methods enhance recommendation robustness in scenarios with limited observations. The causal perspective helps identify which missing interactions represent true negative preferences versus those that are merely unobserved positive preferences, leading to more accurate user modeling.
Third, causal recommendation has advanced objectives beyond accuracy, including explainability (Wachter et al., 2017; Attolou et al., 2024) and fairness (Huang et al., 2022; Zhu et al., 2023). Explainable recommendation benefits from causal reasoning by identifying the actual factors that influence user decisions rather than merely correlative features. Similarly, fairness-aware recommenders use causal techniques to detect and mitigate discrimination against protected groups by modeling the causal relationships between sensitive attributes and recommendation outcomes.
Unlike these previous methods, our approach introduces a novel application of causal inference to sequential recommendation by specifically targeting temporal dynamics. We leverage counterfactual inference to isolate and quantify the causal effect of temporal order on recommendation outcomes, enabling LLMs to better capture the evolution of user preferences over time. By constructing a counterfactual world where temporal information is erased while preserving item content, we can directly measure temporal sensitivity and design specialized tuning objectives that enhance the model’s awareness of both absolute and relative temporal patterns in user interaction sequences.
@@ 3. Preliminaries
[…]
@@ 3.1. Task Formulation
In sequential recommendation scenario, the model predicts the next item a user is likely to interact with based on their historical interaction sequence. Formally, let =$\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ denote the set of items. Given a temporally ordered interaction sequence $\mathcal{S}=(i_{t_{1}},i_{t_{2}},...,i_{t_{|\mathcal{S}|}})$, where $i_{t_{n}}\in\mathcal{I}$ denotes the item interacted with at time $t_{n}$, the goal of the model is to predict the next item $i_{t_{n}+1}$.
@@ 4.1. Causal Analysis
This causal graph provides several important insights. Previous studies have primarily focused on the path $H\rightarrow Y$ or path $I\rightarrow H\rightarrow Y$, corresponding to paradigm adaptation methods and multi-modal integration methods, respectively. The former treats recommendation as a sequence-to-item generation task, while the latter enriches item representations with specialized embeddings. However, the influence of temporal order—captured by the path $T\rightarrow H\rightarrow Y$-has been largely overlooked despite its crucial role in shaping user preferences and interaction patterns.
@@ 4.2. Modeling Temporal Order
To effectively encode temporal information in user interaction sequences, we introduce temporal embedding—a cross-token level position embedding approach that preserves the semantic integrity of each item while explicitly representing its position in the user’s interaction history. Specifically, for sequence $\mathcal{S}$, we assign each item $i_{k}$ a position embedding $p_{k}$ based on its position $k$ in interaction sequence, and add $p_{k}$ to every token of that item. Crucially, rather than applying this embedding at the token level (as in traditional Transformer models), we add the same position embedding $p_{k}$ to every token that constitutes the item. This approach treats each item as a cohesive temporal unit, reinforcing its semantic boundaries within the sequence. We also retain the original token-level positional embeddings to preserve the model’s ability to understand contextual relationships within the prompt. Consequently, the final input embedding for the t-th token is given by:
@@
where $x_{t}$ and $p_{t}$ denote the token embedding and position embedding of the $t$-th token respectively, and $p_{k}$ represents the temporal embedding of the $k$-th item. Besides, $p_{k}$ is set to null if the $t$-th token does not belong to any item. For example, suppose that the movie Shawshank Redemption appears in the third position of $\mathcal{S}$, tokenized into 5 tokens (i.e., ‘Sh’, ‘aw’, ‘sh’, ‘ank’, ‘Redemption’), then each of these 5 tokens will be augmented by the embedding $p_{3}$, signifying that they collectively represent an item that appeared third in the user’s interaction history. This consistent application of temporal embeddings across all tokens of an item ensures that the model perceives the item as a unified entity within the temporal sequence.
This temporal embedding approach offers two primary advantages over standard position embeddings: 1) Item-level semantic coherence: By applying the same temporal embedding to all tokens of an item, we enable the model to treat each item as a cohesive unit during representation learning and inference. This alignment between semantic boundaries and temporal encoding helps the model better leverage item-level prior knowledge and understand items as integral units in the user’s preference evolution. 2)Explicit temporal encoding: By assigning incrementally increasing position embeddings ($p_{1},p_{2},p_{3},\ldots$) to items based on their interaction order, we explicitly encode both absolute and relative temporal relationships. This makes the temporal structure of the interaction sequence more salient to the model, strengthening its ability to detect patterns like recent interest shifts or consistent preferences over time.
@@ 4.5. Discussion
Dataset MovieLens Steam
#Sequences 192,483 151,056
#Items 3,952 3,581
#Interactions 999,611 239,796
@@ 5. Experiment
  • RQ2: How does different position embedding emthods applied in temporal embedding perform?
@@
Table 3. Performance of baselines and CETRec on MovieLens and Steam datasets.
MovieLens Steam
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
Caser 0.0151 0.0126 0.0310 0.0172 0.0355 0.0315 0.0529 0.0367
GRU4Rec 0.0335 0.0279 0.0511 0.0329 0.0397 0.0362 0.0538 0.0402
SASRec 0.0470 0.0265 0.0797 0.0371 0.0544 0.0376 0.0653 0.0411
P5 0.0538 0.0334 0.0911 0.0452 0.0688 0.0468 0.1049 0.0583
BIGRec 0.0985 0.0614 0.1093 0.0650 0.0595 0.0408 0.0687 0.0438
CFT 0.1200 0.0733 0.1299 0.0765 0.0737 0.0498 0.0838 0.0530
CETRec (SinPE) 0.1101 0.0693 0.1209 0.0728 0.0670 0.0436 0.0779 0.0471
CETRec (RoPE) 0.1283 0.0779 0.1365 0.0806 0.0888 0.0635 0.0997 0.0670
@@ 5.1.1. Datasets
We conduct experiments on two real real-wrold datasets: MovieLens [^1] https://grouplens.org/datasets/movielens/ and Steam (Kang and McAuley, 2018). MovieLens dataset is a well-known movie recommendation dataset containing the movies titles, user ratings and timestamps. Steam dataset is a video-games recommendation dataset that records the title of games users have played on the Steam Store. In both datasets, user interactions can be organized into sequences in chronological order.
@@ 5.1.3. Baselines
  • GRU4Rec (Hidasi et al., 2016). An RNN-based recommender that uses Gated Recurrent Units (GRUs) to encode user interaction histories. It effectively captures both short- and long-term dependencies in sequential click data.
  • Caser (Tang and Wang, 2018). A CNN-based model that represents the last $L$ interactions as an $L\times d$ embedding “image” and applies horizontal and vertical convolutional filters. It learns both local item transition patterns and global user preferences for next-item prediction.
  • SASRec (Kang and McAuley, 2018). A Transformer-based recommender that employs masked multi-head self-attention to attend to all previous items in a sequence. It excels at modeling complex dependencies without relying on recurrent connections.
  • P5 (Geng et al., 2022). A T5-based framework that unifies multiple recommendation tasks into a natural-language generation paradigm. P5 maps each task to a specific text prompt, feeds it into T5, and decodes item IDs as generated tokens. By leveraging T5’s pre-trained language understanding, P5 can flexibly handle diverse recommendation subtasks, though it relies on ID tokens for item representation.
  • BIGRec (Bao et al., 2023a). A LLM-based approach for recommendation that converts a user’s interaction history sequence into a textual prompt. The model generates a natural‐language continuation, and a grounding module maps the generated text back to concrete item.
  • CFT (Zhang et al., 2024b). A method that enhances modeling of user behavior sequences through counterfactual fine-tuning. While CFT improves overall sequence understanding, it does not explicitly encode the temporal ordering of interactions.

### v2

# Enhancing Temporal Sensitivity of Large Language Model for Recommendation with Counterfactual Tuning

@@ Abstract.
To address this critical gap, we propose Counterfactual Enhanced Temporal Framework for LLM‑Based Recommendation (CETRec). CETRec is grounded in causal inference principles, which allow it to isolate and measure the specific impact of temporal information on recommendation outcomes. By conceptualizing temporal order as an independent causal factor distinct from item content, we can quantify its unique contribution through counterfactual reasoning—comparing what recommendations would be made with and without temporal information while keeping all other factors constant. This causal framing enables CETRec to design a novel counterfactual tuning objective that directly optimizes the model’s temporal sensitivity, teaching LLMs to recognize both absolute timestamps and relative ordering patterns in user histories. Combined with our counterfactual tuning task derived from causal analysis, CETRec effectively enhances LLMs’ awareness of both absolute order (how recently items were interacted with) and relative order (the sequential relationships between items). Extensive experiments on real-world datasets demonstrate the effectiveness of our CETRec. Our code is available at https://anonymous.4open.science/r/CETRec-B9CE/.
@@ 1. Introduction
In the era of information overload, recommender systems have become essential tools that help users navigate vast amounts of online content (Fan et al., 2023). Sequential recommendation, which models users’ temporally ordered interaction histories to capture evolving preferences, has received increasing attention for producing more personalized and accurate suggestions (Wu et al., 2023; Tang and Wang, 2018; Hidasi et al., 2016; Kang and McAuley, 2018; Rajput et al., 2023). More recently, the rapid advancement of large language models (LLMs) has motivated many works to integrate these models into sequential recommendation systems, leveraging their rich pre-trained knowledge and reasoning capabilities (Bao et al., 2023a; Zheng et al., 2024; Liao et al., 2024; Geng et al., 2022; Bao et al., 2023b; Wei et al., 2024). Temporal information — the chronological order and timing of user interactions — remains central to sequential recommendation, since it directly reflects the evolution of user interests and preference dynamics.
@@
Existing LLM-based recommendation approaches can be largely grouped into two main streams: 1) Paradigm adaptation methods reformulates recommendation as natural language generation (e.g., text completion or response generation) using prompting or fine-tuning to align LLMs’ knowledge with recommendation goals (Bao et al., 2023a; Geng et al., 2022). 2) Multimodal integration represents items as a distinct modality and aligning these embeddings with LLMs’ text representations via cross-modal attention or embedding alignment (Yang et al., 2023; Lin et al., 2023). Despite these advancements, these approaches primarily focus on modeling the semantic aspects of interacted items or optimizing the overall recommendation task, and do not explicitly capture the temporal structure of user interaction sequences. This temporal oversight leaves a significant gap in their ability to model users’ evolving interests and time-dependent preference shifts.Several studies have exposed this weakness: TIN (Zhou et al., 2024) analyzed the correlation between item position and recommendation accuracy, LLM-SRec (Kim et al., 2025) perturbed interaction orders and observed minor performance drops compared to traditional models, these findings collectively suggest that LLM-based recommenders have limited temporal awareness despite diverse modeling strategies.
@@
  • We reveal that existing LLM-based sequential recommenders insufficiently exploit temporal information, due to Transformer architecture limitations and position embeddings designed for NLP.
@@ 2.1. Sequential Recommendation
  • Paradigm Adaption. These methods leverage the reasoning capabilities of LLMs by aligning them with the recommendation paradigm—reformulated as natural language generation tasks—through techniques such as fine-tuning (Geng et al., 2022; Bao et al., 2023a, b; Zhang et al., 2024a), prompting (Dai et al., 2023; Chu et al., 2024), or in-context learning (ICL) (Hou et al., 2024).
Several recent studies have recognized the insufficient utilization of temporal information in interaction sequences by existing recommender systems. TIN (Zhou et al., 2024) examined the correlation between item positions in the sequence and the recommendation results, while LLM-SRec (Kim et al., 2025) evaluated model performance after shuffling the order of interacted items. Both studies revealed clear deficiencies in existing models. To address this issue, TIN adopts Target-aware Temporal Encoding (TTE), integrating it with item ID embeddings to capture semantic–temporal correlations. LLM-SRec distills item embeddings from traditional sequential recommenders, thereby injecting temporal information into LLMs. Tempura (Chu et al., 2024) designs a prompting framework to simulate human reasoning processes, aiming to enhance the temporal awareness of LLMs.
Building on these observations, CETRec injects unified item-level temporal embeddings directly at the LLM input, rather than distributing temporal signals across token-level embeddings or relying solely on distilled representations. This design preserves each item’s lexical semantics while explicitly encoding both absolute and relative temporal relationships, thereby improving the model’s temporal sensitivity and yielding more accurate, time-aware recommendations
@@ 2.2. Causal Inference for Recommendation
Non-causal recommenders typically capture correlations only, which can produce spurious relationships and poor generalization; causal inference instead seeks the underlying data-generating mechanisms, yielding more robust and interpretable recommendations. Prior causal work has concentrated on three practical directions: debiasing (modeling popularity, position/click, and other exposure biases as confounders and applying interventions to correct them) (Zhu et al., 2020; Christakopoulou et al., 2020; Craswell et al., 2008; Wang et al., 2021), using structural causal models to mitigate sparsity and noise by separating genuine preferences from random or missing interactions (Xiong et al., 2021; Zhang et al., 2021; Liu et al., 2022), and moving beyond accuracy toward explainability and fairness by identifying causal drivers of user behavior and detecting/mitigating discriminatory effects (Wachter et al., 2017; Attolou et al., 2024; Huang et al., 2022; Zhu et al., 2023).
Unlike these prior efforts, our work brings causal inference into sequential recommendation with an explicit focus on temporal dynamics: we use counterfactual inference to isolate the causal effect of item order on recommendation outcomes, construct counterfactual worlds that erase temporal signals while preserving item content to measure temporal sensitivity, and introduce tuning objectives that improve LLMs’ modeling of both absolute and relative temporal patterns in user interaction sequences.
@@ 3. Preliminaries
[2pt/2pt] Task Input: The user has played the following video games before: <His_Seq>
@@ 3.1. Task Formulation
In sequential recommendation scenario, the model predicts the next item a user is likely to interact with based on their historical interaction sequence. Formally, let $\mathcal{I}=\{i_{1},i_{2},...,i_{|\mathcal{I}|}\}$ denote the set of items. Given a temporally ordered interaction sequence $\mathcal{S}=(i_{t_{1}},i_{t_{2}},...,i_{t_{|\mathcal{S}|}})$, where $i_{t_{n}}\in\mathcal{I}$ denotes the item interacted with at time $t_{n}$, the goal of the model is to predict the next item $i_{t_{n}+1}$.
@@ 4.1. Causal Analysis
This causal graph provides several important insights. Previous studies have primarily focused on the path $H\rightarrow Y$ or path $I\rightarrow H\rightarrow Y$, corresponding to paradigm adaptation methods and multi-modal integration methods, respectively. The former treats recommendation as a sequence-to-item generation task, while the latter enriches item representations with specialized embeddings. However, the influence of temporal order—captured by the path $T\rightarrow H\rightarrow Y$—has been largely overlooked despite its crucial role in shaping user preferences and interaction patterns.
@@ 4.2. Modeling Temporal Order
4.2.1. Temporal Embedding
To effectively encode temporal information in user interaction sequences, we introduce temporal embedding—a cross-token level position embedding approach that preserves the semantic integrity of each item while explicitly representing its position in the user’s interaction history. Specifically, for sequence $\mathcal{S}$, we assign each item $i_{k}$ a position embedding $p_{k}$ based on its relative position $k$ (i.e., its distance to the most recent item)in interaction sequence, and add $p_{k}$ to every token of that item. Crucially, rather than applying this embedding at the token level (as in traditional Transformer models), we add the same position embedding $p_{k}$ to every token that constitutes the item. This approach treats each item as a cohesive temporal unit, reinforcing its semantic boundaries within the sequence. We also retain the original token-level positional embeddings to preserve the model’s ability to understand contextual relationships within the prompt. Consequently, the final input embedding for the t-th token is:
@@ 4.2.1. Temporal Embedding
where $x_{t}$ and $p_{t}$ denote the token embedding and position embedding of the $t$-th token respectively, and $p_{k}$ represents the temporal embedding of the $k$-th item. Besides, $p_{k}$ is set to null if the $t$-th token does not belong to any item. For example, suppose that the movie Shawshank Redemption appears in the third position of $\mathcal{S}$, tokenized into 5 tokens (i.e., ‘Sh’, ‘aw’, ‘sh’, ‘ank’, ‘Redemption’), then each of these 5 tokens will be augmented by the embedding $p_{3}$, signifying that they collectively represent an item that appeared third in the user’s interaction history.
This temporal embedding approach offers two primary advantages over standard position embeddings: 1) Item-level semantic coherence: By applying the same temporal embedding to all tokens of an item, we enable the model to treat each item as a cohesive unit during representation learning and inference. This alignment between semantic boundaries and temporal encoding helps the model better leverage item-level prior knowledge and understand items as integral units in the user’s preference evolution. 2)Explicit temporal encoding: By assigning incrementally increasing temporal embeddings to items based on their relative interaction order ($p_{9},\ldots,p_{2},p_{1},p_{0}$), we explicitly encode both absolute and relative temporal relationships, allowing later items to have a greater influence on the prediction. This makes the temporal structure of the interaction sequence more salient to the model, strengthening its ability to detect patterns like recent interest shifts or consistent preferences over time.
4.2.2. Connection to Existing Methods
The idea of incorporating temporal embeddings has appeared in several existing methods. In terms of implementation, however, temporal embeddings differ from other approaches. For example, the Target-aware Temporal Encoding (TTE) in TIN adds temporal information to the ID embedding separately within a Transformer-like architecture. This design is essentially similar to conventional positional embeddings. In contrast, our temporal embedding is uniformly added to each item at the input layer of the LLM, providing richer semantic information.
The whole-word embedding in P5 (Geng et al., 2022) also applies additional embeddings to all tokens of an item. However, since it is applied across all tokens in the input sequence, its contribution to modeling item-level temporal information is diluted. Moreover, its encoding order increases sequentially with the token positions, just like standard positional embeddings. This means that items appearing later in the sequence undergo greater changes, thereby weakening their overall influence on the final prediction.
@@ 4.5. Discussion
Dataset MovieLens Steam LastFM
#Sequences 192,483 151,056 46,897
#Items 3,952 3,581 4,606
#Interactions 999,611 239,796 73,510
@@ 5. Experiment
  • RQ2: How does different position embedding methods applied in temporal embedding perform?
@@
Table 3. Performance of baselines and CETRec on MovieLens, Steam, and LastFM datasets.
MovieLens Steam LastFM
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
Caser 0.0667 0.0399 0.1217 0.0573 0.0307 0.0200 0.0488 0.0258 0.0222 0.0145 0.0324 0.0177
GRU4Rec 0.0475 0.0283 0.0900 0.0417 0.0301 0.0198 0.0479 0.0255 0.0130 0.0079 0.0204 0.0103
SASRec 0.0542 0.0322 0.0917 0.0441 0.0269 0.0171 0.0450 0.0229 0.0194 0.0097 0.0352 0.0146
P5 0.0538 0.0334 0.0911 0.0452 0.0688 0.0468 0.1049 0.0583 0.0173 0.0138 0.0273 0.0171
BIGRec 0.0985 0.0614 0.1093 0.0650 0.0778 0.0530 0.0979 0.0596 0.0264 0.0192 0.0356 0.0221
E4SRec 0.1050 0.0660 0.1170 0.0700 0.0810 0.0555 0.0982 0.0615 0.0500 0.0325 0.0600 0.0355
LLaRA 0.1120 0.0710 0.1250 0.0750 0.0850 0.0580 0.0985 0.0635 0.0745 0.0450 0.0880 0.0490
CFT 0.1043 0.0623 0.1241 0.0691 0.0879 0.0606 0.1013 0.0650 0.0711 0.0424 0.0838 0.0467
CETRec (SinPE) 0.1101 0.0693 0.1209 0.0728 0.0670 0.0436 0.0779 0.0471 0.0780 0.0520 0.0910 0.0550
CETRec (RoPE) 0.1283 0.0779 0.1365 0.0806 0.0921 0.0641 0.1072 0.0690 0.0866 0.0591 0.1030 0.0646
@@ 5.1.1. Datasets
We conduct experiments on three real real-wrold datasets: MovieLens [^1] https://grouplens.org/datasets/movielens/ , Steam (Kang and McAuley, 2018) and LastFM (Cantador et al., 2011). MovieLens dataset is a well-known movie recommendation dataset containing the movies titles, user ratings and timestamps. Steam dataset is a video-games recommendation dataset that records the title of games users have played on the Steam Store. LastFM dataset contains the name of artists users listen to collected from the Last.fm music platform. In both datasets, user interactions can be organized into sequences in chronological order.
@@ 5.1.3. Baselines
  • GRU4Rec (Hidasi et al., 2016). GRU4Rec is an RNN-based recommender that uses Gated Recurrent Units (GRUs) to encode user interaction histories.
  • Caser (Tang and Wang, 2018). Caser is a CNN-based model that represents the last $L$ interactions as an $L\times d$ embedding “image” and applies horizontal and vertical convolutional filters.
  • SASRec (Kang and McAuley, 2018). SASRec is Transformer-based recommender that employs masked multi-head self-attention to attend to all previous items in a sequence.
  • P5 (Geng et al., 2022). 5 unifies multiple recommendation tasks into a natural-language generation paradigm, representing items as textual IDs and casting recommendation as conditional text generation.
  • BIGRec (Bao et al., 2023a). BIGRec converts a user’s interaction history sequence into a textual prompt. The model generates a natural‐language continuation, and a grounding module maps the generated text back to concrete item.
  • E4SRec (Li et al., 2023). E4SRec projects item IDs into LLM input embeddings and maps the LLM outputs back to the item space.
  • LLaRA (Liao et al., 2024). LLaRA combines embeddings learned by traditional sequential recommenders and textual expressions of items as prompt to generate the next prediction.
  • CFT (Zhang et al., 2024b). CFT is a method that enhances modeling of user behavior sequences through counterfactual fine-tuning. While CFT improves overall sequence understanding, it does not explicitly encode the temporal ordering of interactions.